OpenStudy (anonymous):

solve the equation -6+3/5x=15

OpenStudy (anonymous):

is it (-6 +3)/5x = 15 or -6 + 3/5x = 15

OpenStudy (anonymous):

-6 + 3/5x = 15

OpenStudy (anonymous):

ORRRR is it -6 + (3/5)x = 15

OpenStudy (anonymous):

its -6 + 3/5x = 15

OpenStudy (anonymous):

-6 + 3/5x = 15 -30x + 3 = 75x 3 = 105x x = 3/105 = 1/35

OpenStudy (anonymous):

\[-6+\frac{3}{5}x=15\] \[-6+\frac{3}{5x}=15\] which one?

OpenStudy (anonymous):

the top one

OpenStudy (anonymous):

lol -30 +3x = 75 3x = 105 x = 35

OpenStudy (anonymous):

good one iant by the way

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

thank you